Find all of the zeros of the function f(x)=6x^3-76X^2+256x-256
Answer by stanbon(75887)

Find all of the zeros of the function f(x)=6x^3-76X^2+256x-256
----
2(3x^3-38x^2 + 128x-128) = 0
I graphed it and found zeroes at x = 2,8/3,and 8
================================
2(3x-8)(x-8)(x-2) = 0
Cheers,
Stan H.